OpenStudy (anonymous):

multiply and simplify : x-6/x^2+4x-32*x-4/x+2

OpenStudy (kropot72):

The denominator of the first term can be factorised as follows:\[x ^{2}+4x-32=(x+8)(x-4)\] The multiplication of the two terms gives: \[(x-6)\div((x+8)(x+2)=(x-6)\div(x ^{2}+10x+16)\]

OpenStudy (anonymous):

Thanks :D